### v1

# Cost Functions in Economic Complexity

Abstract
Economic complexity algorithms, including the Economic Complexity Index (ECI) and Non-Homogeneous Economic Fitness and Complexity (NHEFC), have proven effective at capturing the intricate dynamics of economic systems. Here, we present a fundamental reinterpretation of these algorithms by reformulating them as optimization problems that minimize specific cost functions. We show that ECI computation is equivalent to finding eigenvectors of the network’s transition matrix by minimizing the quadratic form associated with the network’s Laplacian. For NHEFC, we derive a novel cost function that exploits the algorithm’s intrinsic logarithmic structure and clarifies the role of its regularization parameter. Additionally, we establish the uniqueness of the NHEFC solution, providing theoretical foundations for its application. This optimization-based reformulation bridges economic complexity and established frameworks in spectral theory, network science, and optimization. The theoretical insights translate into practical computational advantages: we introduce a conservative, gradient-based update rule that substantially accelerates algorithmic convergence. Beyond advancing our theoretical understanding of economic complexity indicators, this work opens new pathways for algorithmic improvements and extends applicability to general network structures beyond traditional bipartite economic networks.
Keywords Complex Networks  $\cdot$ Economic Complexity Index  $\cdot$ Economic Fitness and Complexity  $\cdot$ Cost Functions
1 Introduction
Economic complexity measures provide valuable insights into the productive capabilities of countries and the structure of international trade [1, 2]. Two of the most widely used measures, the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC), have been instrumental in quantifying the economic diversification of nations and the ubiquity of products [3]. The foundational works by Hidalgo and Hausmann [4, 5] introduced ECI as a method to infer a country’s economic capabilities by analyzing the structure of the country-product network. Subsequently, Tacchella et al. [6, 7] proposed an alternative approach with the Economic Fitness and Complexity method, introducing a non-linear iterative framework to better capture the interplay between country diversification and product competitiveness [8]. Later refinements of EFC, such as the regularized Non-Homogeneous Economic Fitness and Complexity (NHEFC) method [9], aimed to ensure robust convergence and improve its applicability to empirical data.
While these algorithms have demonstrated strong empirical relevance in understanding economic growth and trade networks, their mathematical foundations remain an active area of research. Previous studies [10, 11] have explored the connections between economic complexity indices and spectral methods in network science. For example, in [12] the authors show how ECI and EFC can be integrated into a unified measure incorporating information from both approaches, highlighting their structural similarities despite differences in algorithmic form. In [13, 14], the authors investigate the relation between economic complexity indices and standard notions of centrality, showing that these algorithms can provide valuable insight into the spectral properties of the underlying graph. More recently, ECI and EFC have also been shown to be formally equivalent to network centrality measures [15]. Extending these algorithms to arbitrary network structures, beyond the canonical bipartite formulation, has significantly broadened their applicability and interpretability. Therefore, a deeper theoretical understanding of these approaches is crucial to expanding their potential applications, as already demonstrated in fields such as ecology [16, 17, 18].
Building upon these developments, recent studies have proposed a connection between the EFC algorithm and optimal transport theory [19], particularly in relation to the Sinkhorn-Knopp algorithm [20], which is used to compute doubly stochastic matrices. In this context, economic complexity indices can be interpreted as solutions to optimization problems analogous to those found in optimal transport [21, 22], where cost functions play a central role in defining the equilibrium distributions [23]. Notably, the iterative structure of the NHEFC algorithm resembles the scaling procedures used in transport problems, and its associated cost function exhibits a formal analogy with Kantorovich potentials [24], which characterize the optimal allocation of resources over networks. More generally, the interpretation of economic complexity algorithms as optimization problems [25, 26, 27] provides a unified framework that clarifies their mathematical structure, reveals properties of their solutions, and facilitates connections with other domains.
This study demonstrates that both ECI and NHEFC can be reformulated as optimization problems, each minimizing a distinct cost function. Specifically, we show that the ECI algorithm minimizes an energy function derived from the network’s Laplacian matrix. In contrast, we prove that the cost function underlying NHEFC is strictly convex, thus providing a rigorous foundation for its optimization-based formulation. We also establish convergence to the fixed point—corresponding to the minimum of the cost function—in the case of regular graphs, and prove local convergence for general undirected and unweighted graphs. In addition, we propose a modification to the iterative algorithm that directly follows the gradient descent of the potential, significantly accelerating numerical convergence.
The results presented in this paper enhance the interpretability of economic complexity algorithms and pave the way for methodological improvements and new applications in network-based data analysis. The reformulation of these indices as cost-minimization problems contributes to a more rigorous theoretical understanding of economic complexity and opens new perspectives for their application beyond traditional bipartite economic networks.
2 Materials and Methods
2.1 Data
For the analysis presented in this study, we primarily rely on synthetic regular graphs with a variable number of nodes $N$ and fixed degree $K$, identical for all nodes. These graphs are artificially generated to match the structural properties required for the analytical investigations developed in the following sections. Since many of our theoretical results are derived under the assumption of regularity, these synthetic graphs serve as a controlled setting to validate and illustrate the analytical findings.
We also consider the well-known Zachary Karate Club network to assess the algorithm’s behavior on real-world data. This dataset, originally collected by Wayne Zachary in 1977 [28, 29], represents the social interactions among members of a university karate club. Each node corresponds to a club member, while edges represent observed social ties between them. The network is undirected and is commonly used as a benchmark for community detection algorithms, particularly for identifying the two factions that formed due to a conflict between the club’s instructors.
2.2 Previous work
In our previous work [15, 23], we described two complementary methods for quantifying economic complexity: the Economic Complexity Index (ECI) and the Economic Fitness and Complexity (EFC). Both approaches are based on the country–product binary matrix $\mathbf{M}$. In this matrix, the entry $M_{cp}$ is set to 1 if country $c$ significantly exports product $p$, and 0 otherwise, as determined by the Revealed Comparative Advantage (RCA) criterion [30].
For ECI, one defines the diversification of a country and the ubiquity of a product as
$k_{c}=\sum_{p=1}^{N_{p}}M_{cp}\quad\text{and}\quad k_{p}=\sum_{c=1}^{N_{c}}M_{cp},$
and updates the scores iteratively via the method of reflections [4]:
$F^{(n)}_{c}=\frac{1}{k_{c}}\sum_{p=1}^{N_{p}}M_{cp}Q^{(n-1)}_{p},\quad Q^{(n)}_{p}=\frac{1}{k_{p}}\sum_{c=1}^{N_{c}}M_{cp}F^{(n-1)}_{c},$
with initial conditions $F^{(0)}_{c}=k_{c}$ and $Q^{(0)}_{p}=k_{p}$.
We can recast ECI in a spectral framework by writing the iterative process in matrix form using the diagonal matrices $\mathbf{D}$ and $\mathbf{U}$ (with elements $k_{c}$ and $k_{p}$, respectively), so that
$\vec{F}^{(n)}=\mathbf{D}^{-1}\mathbf{M}\,\vec{Q}^{(n-1)}~{}~{}\text{and}~{}~{}\vec{Q}^{(n)}=\mathbf{U}^{-1}\mathbf{M}^{T}\,\vec{F}^{(n-1)}.$
After two iterations, these lead to the transition matrices $\mathbf{N}_{1}=\mathbf{D}^{-1}\mathbf{M}\mathbf{U}^{-1}\mathbf{M}^{T}$ and $\mathbf{N}_{2}=\mathbf{U}^{-1}\mathbf{M}^{T}\mathbf{D}^{-1}\mathbf{M}$ whose second eigenvectors (beyond the trivial constant principal eigenvector) capture the nontrivial complexity rankings [10, 11].
In contrast, EFC employs a non-linear map [6]. To ensure robust convergence, we used the regularized Non-Homogeneous EFC (NHEFC) given by [9]
$\begin{cases}{{F}}^{(n)}_{c}=\delta+\sum_{p=1}^{N_{p}}M_{cp}/{{P}}^{(n-1)}_{p}\\ {{P}}^{(n)}_{p}=\delta+\sum_{c=1}^{N_{c}}M_{cp}/{{F}}^{(n-1)}_{c}\end{cases}$ (1)
where $F^{(0)}_{c}=P^{(0)}_{p}=1$ and $\delta=10^{-3}$.
By introducing the adjacency matrix $\mathbf{A}$ and the diagonal matrix $\mathbf{K}$ with $k_{c}$ and $k_{p}$ on its diagonal, we can provide a mono-partite representation that recasts both ECI and NHEFC as iterative maps over a unified vector $\vec{V}$ of dimension $N=N_{c}+N_{p}$.
In particular, we write ECI as
$\vec{V}^{(n)}=\mathbf{K}^{-1}\mathbf{A}\;\vec{V}^{(n-1)}~{}~{}\mbox{with}~{}~{}~{}\vec{V}^{(0)}=\vec{k}.$ (2)
We similarly write NHEFC as
$\vec{V}^{(n)}=\delta\cdot\vec{1}+\mathbf{A}\cdot(\vec{V}^{(n-1)})^{-1}~{}~{}\mbox{with}~{}~{}~{}\vec{V}^{(0)}=\vec{1},$ (3)
or explicitly
$V^{(n)}_{i}=\delta+\sum_{j=1}^{N}\frac{A_{ij}}{V^{(n-1)}_{j}}~{}~{}\mbox{with}~{}~{}~{}V_{i}^{(0)}=1,$ (4)
thereby generalizing the approach, originally designed for bipartite structures, to any simple-graph structure. In Eq. (3), the notation $(\vec{V})^{-1}$ stands for taking the reciprocal of each element of the vector $\vec{V}$. This mono-partite formulation is a key generalization, allowing the ECI and NHEFC concepts to be applied to any arbitrary network structure, not just the bipartite country-product networks for which they were originally designed.
2.3 Cost Functions
In this section, we show that both ECI and NHEFC can be understood as optimization algorithms, as they can be reformulated as problems of minimizing specific cost functions. In many optimization settings, the cost (or objective) function is defined from the outset, and iterative procedures are devised to find its minimum. In our case, the process is reversed: starting from the iterative maps in Eq. (2) and Eq. (3), we reconstruct the corresponding cost functions by integration. The cost function associated with ECI relates directly to the standard eigenvector problem and can be interpreted within that framework. For NHEFC, the connection to optimization has been previously noted due to structural similarities between the fitness algorithm and the Sinkhorn-Knopp algorithm from optimal transport theory [31]. However, while prior work has identified this formal analogy [19], our study provides the first explicit derivation of NHEFC’s cost function. This explicit formulation enables us to prove the function’s convexity rigorously and subsequently develop a more efficient gradient-based algorithm. In the following, we derive the explicit expression of the cost function associated with the NHEFC algorithm. Full details of the derivations for both ECI and NHEFC are provided in Appendix A.
2.3.1 ECI
By integrating Eq. (2), we obtain that the solution of the algorithm is equivalent to finding the minimum of the following cost function:
$U(\vec{V})=-\dfrac{1}{2}\sum_{i,j}A_{ij}V_{i}V_{j}+\dfrac{1}{2}\sum_{i}k_{i}V_{i}^{2}.$ (5)
This function has the same form as the cost function used in standard eigenvector-finding algorithms [32]. It is also formally related to a classical harmonic potential [33] or spin-system Hamiltonian, where the first term represents a standard quadratic interaction, and the second acts as a regularization.
We can also express the cost function Eq. (5) in terms of the Laplacian matrix $L_{ij}=k_{i}\delta_{ij}-A_{ij}$ associated with the adjacency matrix $A_{ij}$:
$U(\vec{V})=\frac{1}{4}\sum_{i,j}A_{ij}(V_{i}-V_{j})^{2}=\dfrac{1}{2}\sum_{i,j}L_{ij}V_{i}V_{j}.$ (6)
The minimum of this function occurs when $V_{i}=V_{j}$ for all $i$ and $j$, which corresponds to the constant vector. This is also the first eigenvector of the Laplacian matrix, representing the solution that the ECI algorithm approaches after many iterations.
2.3.2 NHEFC
In this case, integrating Eq. (4) gives the following cost function:
$U(\vec{V})=\dfrac{1}{2}\sum_{i,j}A_{ij}\dfrac{1}{V_{i}V_{j}}-\sum_{i}\log\dfrac{1}{V_{i}}+\delta\sum_{i}\dfrac{1}{V_{i}}.$ (7)
The cost function Eq. (5) associated with ECI resembles a standard harmonic potential, where the quadratic term promotes positive interactions (i.e., attraction) between connected nodes. In contrast, the cost function of EFC features interaction terms that involve the inverse of the variables. This results in a locally repulsive dynamics, which promotes differentiation between neighboring nodes, opposite to the attractive mechanisms typical of harmonic potentials [33]. The expression Eq. (7) also exhibits formal analogies with Kantorovich potentials in optimal transport theory [24].
The repulsive nature of the EFC formulation is already evident in the map Eq. (4), where each fitness value is computed as a weighted average of the inverse of its neighbors. This structure inherently promotes diversity in node values. As a result, and as thoroughly analyzed in [15], fitness centrality is particularly effective in identifying structural vulnerability in networks: it tends to assign high values to nodes connected to weak or fragile ones.
In addition to the interaction term, the cost function includes two regularization terms, both essential for preventing divergence. The first is a logarithmic barrier, which regularizes the contribution of the inverse variables. This term also reveals that the natural scale of the fitness variables is logarithmic, as already pointed out in several cases [6, 19]. However, this contribution alone is insufficient to prevent the divergence caused by the hyperbolic nature of the interaction. The second regularization term, introduced via the constant $\delta$, ensures full algorithm convergence. The parameter $\delta$ can be interpreted as a Lagrange multiplier that enforces the constraint $\sum_{i}V_{i}^{-1}=\text{const}$.
3 Results
In this section, we analyze the convergence properties of ECI and NHEFC. For ECI, properties such as the convexity of the potential and the existence of a fixed point follow directly from the fact that ECI is equivalent to computing the second eigenvector of the matrix $\mathbf{K}^{-1}\mathbf{A}$. The following part focuses on the convergence of NHEFC across different network structures, starting from a simple two-node undirected graph and extending to regular graphs and arbitrary topologies. The analysis establishes local convergence criteria near fixed points and explores the influence of the non-homogeneous parameter $\delta$ on the algorithm’s behavior. Finally, we introduce an alternative formulation of the fitness algorithm that follows the gradient of the potential function. This conservative formulation defines the most direct path to the fixed point and significantly reduces the number of iterations required for convergence.
3.1 Convergence properties of ECI
As previously noted, computing ECI values is equivalent to extracting the eigenvectors of the matrix $\mathbf{K}^{-1}\mathbf{A}$, as defined by the map in Eq. (2). Since $\mathbf{K}^{-1}\mathbf{A}$ is row-normalized, it acts as a transition matrix. By the Perron–Frobenius theorem, this map converges as $n\to\infty$ to the principal eigenvector of the transition matrix. This eigenvector corresponds to the eigenvalue $\lambda_{1}=1$ and is the constant vector $\vec{V}=(1,\dots,1)$. This result also follows directly from the cost function in Eq. (6), which attains its global minimum when all $V_{i}$ are equal. At this point, the cost is $U(\vec{V})=0$. Moreover, since the cost function in Eq. (6) is a quadratic form defined by the Laplacian network matrix, it is convex over the entire domain. Therefore, the map Eq. (2) has a unique fixed point, and the Perron–Frobenius theorem ensures convergence.
Since this fixed point (the constant vector) does not provide a meaningful solution of the problem of ranking countries or products, the original formulation of ECI [4] prescribes stopping the iteration after a fixed number of steps. As discussed in several works [12, 23, 1], the intermediate steps of the iteration retain information about the subdominant eigenvectors, particularly the second one. As a result, ECI can be obtained in two equivalent ways: (i) by iterating the map Eq. (2) for a fixed number of steps, or (ii) by directly computing the second eigenvector of the transition matrix $\mathbf{K}^{-1}\mathbf{A}$. This eigenvector maximizes the partitioning of the network, in line with the interpretation of ECI as a spectral algorithm for community detection [12].
3.2 Convergence properties of NHEFC
3.2.1 Uniqueness of solution via convexity analysis
Figure 1: Energetic landscape of the fitness algorithm. a) Cost function from Eq. (7) for regular complete graphs with varying number of nodes $N$ and $K=N-1$. The potential $U(V)$ is plotted as a function of a single variable $V$, since by symmetry all node variables can be set equal. The curves for different values of $N$ demonstrate the convexity of the potential. Each curve exhibits a global minimum corresponding to the fixed point in Eq. (11), which for small $\delta$ can be approximated as $V^{*}\approx\sqrt{K}$. b) Cost function around the fixed point $V^{*}$ for the ZKC network [29], plotted as a function of the fitness value $V_{i}$ for specific nodes. Each curve shows how the potential varies around the fixed point $V_{i}^{*}$ for that node. Nodes with higher fitness display larger values of $V_{i}^{*}$, i.e., their minima lie farther to the right. The legend reports the node degrees $d$ for comparison. As shown, the highest fitness nodes contribute most to the total potential, while low-fitness nodes provide only local corrections. Each curve exhibits a unique, global minimum. The curves are convex, though the logarithmic scale on the vertical axis may create a misleading appearance of concavity in some regions.
To establish the uniqueness of the solution to the NHEFC equations, we investigate the properties of the potential function $U(\vec{V})$ defined in Eq. (7). We consider the form of the potential $U(\vec{Z})$ expressed in terms of the natural fitness variables $Z_{i}=-\log V_{i}$, as derived in Appendix A:
$U(\vec{Z})=\dfrac{1}{2}\sum_{i,j}A_{ij}e^{Z_{i}}e^{Z_{j}}-\sum_{i}Z_{i}+\delta\sum_{i}e^{Z_{i}}$ (8)
Specifically, we demonstrate its strict convexity by analyzing its Hessian matrix. The second partial derivatives of $U$ with respect to the variables $Z_{i}$ yield the Hessian matrix elements:
$H_{ij}=\frac{\partial^{2}U}{\partial Z_{i}\partial Z_{j}}=\delta_{ij}e^{Z_{i}}\left(\sum_{k}A_{ik}e^{Z_{k}}+\delta\right)+A_{ij}e^{Z_{i}}e^{Z_{j}},$ (9)
where $\delta_{ij}$ represents the Kronecker delta.
This Hessian matrix allows for a useful interpretation rooted in graph theory. Let us consider an associated weighted graph where the edge weights are given by $W_{ij}=A_{ij}e^{Z_{i}}e^{Z_{j}}$ (non-zero only if $A_{ij}=1$) and the strength of each node $i$ is $S_{i}=\sum_{k}W_{ik}=e^{Z_{i}}\sum_{k}A_{ik}e^{Z_{k}}$. With these definitions, the Hessian matrix $\mathbf{H}$ (with components $H_{ij}$) can be decomposed as:
$\mathbf{H}=\mathbf{Q}+\mathbf{\Delta}.$
Here, $\mathbf{Q}$ is the signless Laplacian matrix of the weighted graph, defined by $Q_{ij}=\delta_{ij}S_{i}+W_{ij}$. From a well-established property in spectral graph theory, the signless Laplacian $\mathbf{Q}$ of any graph (weighted or unweighted) is positive semi-definite; that is, all of its eigenvalues are non-negative ($\lambda(\mathbf{Q})\geq 0$). The second term, $\mathbf{\Delta}$, is a diagonal matrix with entries $\Delta_{ii}=\delta e^{Z_{i}}$ and zeros elsewhere ($\Delta_{ij}=\delta_{ij}\delta e^{Z_{i}}$).
Since the EFC algorithm assumes $\delta>0$ and the exponential function ensures $e^{Z_{i}}>0$ for all $i$, the diagonal entries of $\mathbf{\Delta}$ are strictly positive. Adding a diagonal matrix with strictly positive entries ($\mathbf{\Delta}$) to a positive semi-definite matrix ($\mathbf{Q}$) results in a matrix ($\mathbf{H}$) that is positive definite. This means all eigenvalues of the Hessian matrix $\mathbf{H}$ are strictly positive ($\lambda(\mathbf{H})>0$) for any state $\vec{Z}$.
A fundamental result from multivariate calculus states that if the Hessian matrix of a function is positive definite throughout its domain, the function is strictly convex. Therefore, the potential function $U(\vec{Z})$ is strictly convex. A strictly convex function possesses, at most, one point where its gradient vanishes. This stationary point corresponds precisely to the condition $\dot{Z}_{i}=0$ in the dynamical system formulation (Eq. (19)), which in turn defines the fixed point of the EFC iterative map Eq. (18). Consequently, the strict convexity of $U(\vec{Z})$ guarantees that this fixed point is not only a minimum but the unique global minimum of the potential function, ensuring the uniqueness of the NHEFC solution.
The role of $\delta$ can also be analyzed through its influence on the cost function bounds. From Eq. (4), using $k_{\max}$ as the maximum node degree in the network, we obtain the constraint $\delta\leq V_{i}\leq\delta+k_{\max}/\delta$. In terms of the logarithmic variables $Z_{i}$, this translates to the bounds $-\log(\delta+k_{\max}/\delta)\leq Z_{i}\leq-\log(\delta)$. By incorporating these inequalities into Eq. (20) and retaining only the leading terms, we establish the global bounds on the potential function,
$-N\log(1/\delta)\leq U\leq E/\delta^{2},$ (10)
where $N$ is the number of nodes and $E$ is the number of links in the network. This result demonstrates that $\delta$ effectively regulates the potential function, preventing divergence and ensuring that $U$ remains bounded.
Figure 1 illustrates the shape of the potential function $U(\vec{Z})$ for the EFC algorithm. The left panel shows the case of regular graphs of different sizes, where all nodes are symmetric and share the same fitness value (Eq. (11)). In this case, the potential is plotted as a function of a single variable, confirming its strict convexity. The right panel shows the potential landscape for the Zachary Karate Club network. Here, each node has a distinct fitness value. We visualize the potential $U(\vec{Z})$ level curves by varying one variable at a time around the fixed point. These cross-sections show that the potential is convex along all directions in the fitness space. Moreover, the shape of each curve reveals how much each node contributes to the total potential. Nodes with high fitness (indicated by the location of the minimum in their respective potential sections) induce the steepest variations in the potential. This suggests that central or structurally important nodes contribute more significantly to the system’s energetic configuration, while peripheral nodes have a more negligible effect and contribute only local corrections.
3.2.2 Solution in the case of two connected nodes
Figure 2: Convergence of trajectories in the phase space. a) Bidimensional phase space of variables $V_{1}$ and $V_{2}$ for a complete graph with $N=10$. The colormap shows the value of the potential, with the global minimum corresponding to the fixed point $(V_{1}^{*},V_{2}^{*})\approx(3,3)$ (Eq.(11)). White points indicate several initial conditions, while the lines show the convergence of the original map defined in Eq. (4). Trajectories are shown every ten steps. b) Convergence of several trajectories, starting from different initial conditions, toward the fixed point. Only one variable, $V$, is plotted, since by the symmetry of the graph, all node variables are equivalent. Trajectories are shown every two steps, due to the oscillatory behavior of the map. c) Bidimensional phase space of $V_{1}$ and $V_{2}$ for the ZKC network [29]. The global minimum of the potential corresponds to the fixed point $(V_{1}^{*},V_{2}^{*})\approx(3.2,2.2)\times 10^{3}$. White points indicate various initial conditions, and the lines show the trajectories under Eq. (4), plotted every two steps. d) Convergence of several trajectories to the fixed point for variables $V_{1}$ and $V_{2}$, starting from different initial conditions.
The specific case of a network with only two connected nodes ($N=2$) offers a tractable scenario where the algorithm’s convergence can be analyzed in full detail (refer to Appendix B for a complete treatment). By examining the dynamics over two consecutive steps, it is found that each component $V_{j}$ evolves independently according to the same one-dimensional iterative map: $V^{(n+2)}_{j}=f(V^{(n)}_{j})$, where the function is defined as $f(x)=\delta+x/(1+\delta x)$. For any positive value of the parameter $\delta$, this function $f(x)$ has derivative $10<f^{\prime}(x)<$ 1 for $x>0$, guaranteeing the existence of a unique, globally attracting fixed point within the positive domain. This fixed point is given by:
$x^{*}=\frac{1}{2}\left(\sqrt{4+\delta^{2}}+\delta\right).$
Consequently, for $\delta>0$, the sequence of vectors $\vec{V}^{(n)}$ is guaranteed to converge to the unique equilibrium state $\vec{V}^{*}=(x^{*},x^{*})$ regardless of the positive starting point $\vec{V}^{(0)}$. This fixed point $\vec{V}^{*}$ also corresponds to the unique global minimum of the associated potential function $U(\vec{V})$. The analysis highlights the critical stabilizing role of $\delta$: if $\delta$ is set to zero, the function becomes $f(x)=x$, and convergence is lost; the system typically exhibits persistent period-2 oscillations instead of settling at the fixed point $(1,1)$. Thus, $\delta>0$ acts like a damping mechanism, ensuring asymptotic stability rather than mere stability.
3.2.3 Solution in the case of a regular graph
In the specific case where the underlying network corresponds to a regular graph, meaning every node $i$ has the same degree $K=\sum_{j}A_{ij}$, the convergence properties of the iterative map can be analytically determined (more details are provided in Appendix C). The analysis relies on tracking the minimum, $V^{(n)}_{m}$, and maximum, $V^{(n)}_{M}$, values of the fitness vector components at each iteration $n$. By establishing recursive inequalities, it is shown that the evolution of these bounds is governed by an auxiliary one-dimensional map $X^{(n+1)}=f_{K}(X^{(n)})$. Provided that the parameter $\delta>0$, this map possesses a unique, globally attracting fixed point. This structure rigorously proves that for any positive initial condition $\vec{V}^{(0)}$, the iterative process
$V^{(n)}_{i}=\delta+\sum_{j}\frac{A_{ij}}{V^{(n-1)}_{j}}$
converges exponentially fast to a unique fixed point $\vec{V}^{*}$, where all components are identical and equal to
$V^{*}=\sqrt{K+\tfrac{1}{4}\delta^{2}}+\delta.$ (11)
This fixed point $\vec{V}^{*}$ represents the unique positive solution to the stationarity condition $V_{i}=\delta+\sum_{j}A_{ij}/V_{j}$ and also corresponds to the unique global minimum of the associated cost function $U(\vec{V})$ within the positive orthant. The positivity of $\delta$ is crucial, as the argument for convergence and uniqueness breaks down when $\delta=0$.
The specific shape of the cost function for the EFC algorithm in the case of a regular graph is shown in Fig. 1a. The potential curves highlight the fixed point, corresponding to the function’s global minimum. These curves also clearly demonstrate the strict convexity of the cost function across the entire fitness space. Additionally, Fig. 2a displays the potential landscape in the two-dimensional phase space defined by the first two components of the fitness vector, $V_{1}$ and $V_{2}$, for a regular graph with $N=10$ and degree $K=9$. The figure shows the trajectories of the algorithm starting from various initial conditions, all of which converge to the fixed point, located at $V\approx\sqrt{K}=3$.
3.2.4 Local and general convergence for a general graph
For general graph structures, while global convergence is not guaranteed by the methods used for regular graphs, the local stability of the algorithm’s unique fixed point $\vec{V}^{*}$ can be established using Lyapunov stability theory (more details in Appendix D). The approach utilizes the potential function $U(\vec{V})$, whose unique minimum coincides with $\vec{V}^{*}$, as a Lyapunov function candidate. The analysis focuses on the change in the potential over one iteration, $\Delta U=U(\vec{F}(\vec{V}))-U(\vec{V})$, for states $\vec{V}$ confined to a sufficiently small neighborhood around the fixed point $\vec{V}^{*}$. By performing a Taylor expansion of $\Delta U$ and carefully analyzing its terms, particularly relating the gradient term $\nabla U(\vec{V})\cdot(\vec{F}(\vec{V})-\vec{V})$ to the step size and bounding the second-order (Hessian) term using properties of the fixed point and inequalities like Cauchy-Schwarz, it can be demonstrated that $\Delta U$ is strictly negative for any $\vec{V}\neq\vec{V}^{*}$ within this neighborhood, provided $\delta>0$. To leading order in the displacement $\vec{D}=\vec{F}(\vec{V})-\vec{V}$, the change is found to be approximately
$U(\vec{F}(\vec{V}))-U(\vec{V})\approx-\frac{\delta}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}$
Since this expression is strictly negative for any non-zero displacement $\vec{D}$ (given $\delta>0$), it confirms that $U(\vec{V})$ strictly decreases along the iterative trajectory as it approaches $\vec{V}^{*}$. This rigorously establishes the local asymptotic stability of the unique fixed point $\vec{V}^{*}$ for the dynamics on general graphs when $\delta>0$.
While a general proof of convergence is difficult for arbitrary network structures, this property can be investigated empirically by analyzing the behavior of multiple trajectories initialized from different starting points. We carried out this analysis for the Zachary Karate Club network, as shown in Fig. 2b. The results indicate that, regardless of the initial condition in the phase space, all trajectories converge to the same fixed point. Although a formal proof is lacking, this evidence suggests that global convergence may hold more generally.
3.2.5 Reformulation of the map via gradient descent
Figure 3: Efficient convergence of the conservative map compared to the original map. a) Trajectories of the potential $U(\vec{V})$ (Eq. (7)) for a complete graph with $N=10$. Following the original map, the dynamics exhibit strong period-2 oscillations, indicating algorithm instability. In contrast, the trajectory computed by descending the cost function gradient using the conservative map (Eq. (12)) shows no oscillatory behavior and converges more efficiently to the fixed point. It requires significantly fewer steps to reach a stable solution. b) Trajectories of the potential for the ZKC network. Also in this case, the conservative map converges considerably faster and avoids oscillations. These results suggest that, for practical computations, the original map defined by Eq. (4) should be replaced with the new formulation in Eq. (12).
The reformulation of the fitness algorithm as the minimization of a cost function has practical computational advantages, as shown in Appendix A. When the dynamics are expressed in the logarithmic variables $Z_{i}=-\log V_{i}$, the associated vector field becomes irrotational. This allows us to define a conservative (i.e., gradient-driven) map that follows the steepest descent of the potential function $U(\vec{Z})$, leading to faster convergence. The update rule is:
$Z_{i}^{(n)}=Z_{i}^{(n-1)}+1-\delta e^{Z_{i}^{(n-1)}}-e^{Z_{i}^{(n-1)}}\sum_{j=1}^{N}A_{ij}e^{Z_{j}^{(n-1)}}$ (12)
This iteration differs from the standard map in Eq. (4), as it explicitly follows the cost function gradient in the transformed variables. In Fig. 3, we compare the convergence behavior of both maps. The conservative map exhibits significantly faster convergence and eliminates the oscillatory behavior observed in the original formulation. For practical computations, this gradient-based update rule should be preferred. If necessary, the results can be mapped back to the original variables via $V_{i}=e^{-Z_{i}}$.
4 Discussion
This study establishes the mathematical foundations of two prominent economic complexity measures, ECI and NHEFC, by reformulating them as optimization problems. This approach unifies these algorithms with well-established network science techniques and provides a deeper theoretical understanding and significant practical improvements. Our findings demonstrate that the iterative algorithms for both ECI and NHEFC can be derived as procedures to minimize specific cost functions, revealing the underlying mechanics of each framework.
A key result of our analysis is the contrasting nature of the cost functions associated with ECI and NHEFC. The ECI cost function is a quadratic form related to the network’s Laplacian matrix, analogous to a classical harmonic potential. This formulation promotes attraction, where connected nodes tend to have similar values, which aligns with the original ECI’s established interpretation as a spectral algorithm for network partitioning. In stark contrast, the NHEFC cost function features interaction terms involving the inverse of the node variables, resulting in a locally repulsive dynamic. This repulsive mechanism promotes differentiation among neighboring nodes. This explains the effectiveness of fitness centrality in identifying structural vulnerabilities, as it tends to assign high fitness values to nodes connected to weaker ones. More broadly, it provides a rigorous energetic perspective on the fitness solution, as initially suggested in previous work.
Furthermore, our derivation of the NHEFC cost function clarifies the roles of its components. The function reveals that the natural scale of the fitness variables is logarithmic, a property observed empirically but not formally derived. Crucially, the analysis highlights the role of the regularization parameter $\delta$. This parameter acts as a damping mechanism that ensures the algorithm’s convergence by preventing the divergence caused by the hyperbolic nature of the interactions. Theoretically, the most significant contribution is the proof of the uniqueness of the NHEFC solution. By transforming the potential into its natural logarithmic variables and analyzing its Hessian matrix, we demonstrated that the potential is strictly convex. This strict convexity guarantees a unique global minimum, which corresponds to the algorithm’s fixed point, thus resolving any ambiguity about the uniqueness of the solution.
The optimization framework also offers significant practical and methodological advancements. For NHEFC, we have shown that the dynamics expressed in logarithmic variables are irrotational. This implies that the most efficient computational path to the solution is a direct gradient descent of the potential function. This leads to a new “conservative map” (Eq. (12) in the original text) that converges significantly faster and avoids the period-2 oscillations characteristic of the original map. For practical applications, we recommend that the fitness algorithm be computed using this new formulation.
Interpreting economic complexity algorithms through potential functions opens the way to a novel perspective on their application. In economic networks, the potential function provides an interpretable energetic landscape for each node, capturing both local and global structural forces. This framing of the cost function as a potential opens up a rich conceptual framework. Given the demonstrated strict convexity of the potential, the energetic landscape is characterized by a single, unique global minimum, representing the system’s sole optimal configuration. At any other point, the potential $U(\vec{V})$ quantifies the “stress” or deviation from this ideal state. This picture re-frames economic resilience not as the depth of a potential well, but as the steepness of the landscape around the minimum; a steeper potential implies stronger restorative forces that pull the system back to its optimum after a shock. Consequently, structural transformation is reconceptualized not as a jump between different stable states but as a dynamic change of the landscape itself, where an alteration in the network’s topology reshapes the cost function and shifts the location of the unique minimum. Finally, the geometry of this landscape can inform measures of systemic risk, as a “flatter” potential in certain directions would indicate a system with weak restorative forces, vulnerable to large and inefficient deviations from its optimal configuration. This representation could provide powerful insights into the organization and resilience of the economic fabric, offering a new tool for empirical investigations in regional development, economic geography, and beyond.
Finally, this work broadens the applicability and the theoretical interpretation of economic complexity measures. By reformulating the underlying algorithms as optimization problems, we establish formal connections to broader scientific frameworks, including Kantorovich potentials in optimal transport theory and spectral methods in graph theory through the Laplacian matrix. Crucially, expressing the dynamics in a mono-partite setting removes the constraint of bipartite trade networks, allowing these measures to be applied to arbitrary graph structures. This generalization paves the way for their adoption in diverse contexts, such as ecological systems, infrastructure networks, and spatial economic analysis, significantly extending their scope and relevance.
Author contributions:
V.D.P.S. conceptualized and led the study, including methodology, validation, and initial draft writing. A.B. contributed to methods and software development and created all visualizations. P.B. contributed to the formal analysis. All authors collaborated on the formal analysis and the final manuscript. All authors have read and agreed to the published version of the manuscript.
Funding:
This work was funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation and Technology as part of the project GZ 2021-0.664.668.
Acknowledgments:
We acknowledge valuable discussions with E. Calò, G. De Marzo, D. Mazzilli, and A. Patelli.
Conflicts of interest:
The authors declare no conflict of interest.
Abbreviations:
The following abbreviations are used in this manuscript:
ECI Economic Complexity Index EFC Economic Fitness and Complexity NHEFC Non-Homogeneous Economic Fitness and Complexity RCA Revealed Comparative Advantage
Appendix A Derivation of Cost Functions
Here, we provide the analytical details of the computations for both ECI and NHEFC algorithms. We start with the simplest case, namely ECI.
A.1 ECI
To express the cost function, we must formulate Eq. (2) as a map. This involves subtracting $V_{i}^{(n-1)}$ from both sides of the equations and considering the left-hand side $V_{i}^{(n)}-V_{i}^{(n-1)}\approx\frac{d}{dn}V_{i}=\dot{V_{i}}^{(n-1)}$ as an approximate derivative. The resulting map is expressed as:
$\dot{V_{i}}^{(n-1)}=\dfrac{1}{k_{i}}\sum_{j=1}^{N}A_{ij}V_{j}^{(n-1)}-V_{i}^{(n-1)}$ (13)
Here, the dot represents the derivative with respect to the iteration index (n), and $N=N_{c}+N_{p}$ denotes the total number of nodes in the network, irrespective of their specific nature. The right-hand side of the equation can be interpreted as the vector field driving the dynamics to convergence.
It is convenient to introduce a change of variable to make the network’s adjacency matrix symmetric. This is required to compute the cost function of the map. Introducing the variables $Z_{i}=\sqrt{k_{i}}V_{i}$, Eq. (13) reads as:
$\dot{Z}_{i}=\sum_{j=1}^{N}\dfrac{A_{ij}}{\sqrt{k_{i}k_{j}}}Z_{j}-Z_{i}$ (14)
The potential can now be obtained by integrating this map, but the Schwartz condition must be satisfied. This is equivalent to requiring that the vector field is irrotational (to be more precise, the 2-form derived from the vector field 1-form is zero):
$\dfrac{\partial\dot{Z_{i}}}{\partial Z_{h}}=\dfrac{\partial\dot{Z_{h}}}{\partial Z_{i}}$ (15)
Given that the full adjacency matrix $A_{ij}$ is symmetric by construction, both quantities are equal to $\dfrac{A_{ih}}{\sqrt{k_{i}k_{h}}}+\delta_{ih}$.
This ensures the existence of a potential function in the variables $Z_{i}$ with the property that $-\nabla_{Z_{i}}U=\dot{Z_{i}}$. From this we recover the explicit expression by integrating Eq. (14):
$U(\vec{Z})=-\dfrac{1}{2}\sum_{i,j}\dfrac{A_{ij}}{\sqrt{k_{i}k_{j}}}Z_{i}Z_{j}+\dfrac{1}{2}\sum_{i}Z_{i}^{2}.$ (16)
The factor $1/2$ avoids counting each link twice since the matrix $A_{ij}$ is symmetric. Replacing back the original variables $V_{i}=Z_{i}/\sqrt{k_{i}}$, the cost functions is expressed as:
$U(\vec{V})=-\dfrac{1}{2}\sum_{i,j}A_{ij}V_{i}V_{j}+\dfrac{1}{2}\sum_{i}k_{i}V_{i}^{2}$ (17)
We remark that the interpretation as a potential function is possible only using the natural variables $Z_{i}$. The minimum of Eq. (16) coincides with the minimum of Eq. (17), but the dynamics leading to the minimization may not be equivalent.
A.2 NHEFC
For NHEFC, after removing the isolated nodes the map equivalent to Eq. (13) is:
$\dot{V_{i}}^{(n-1)}=\delta+\sum_{j=1}^{N}\frac{A_{ij}}{V^{(n-1)}_{j}}-V_{i}^{(n-1)}$ (18)
The vector field associated with the system is not irrotational, as Eq. (15) are not satisfied. Through an appropriate change of variable $Z_{i}=-\log V_{i}$, Eq. (18) can be transformed as
$\dot{Z_{i}}=1-\delta e^{Z_{i}}-e^{Z_{i}}\sum_{j=1}^{N}A_{ij}e^{Z_{j}}$ (19)
This map is now irrotational, and the change of variable naturally reveals the intrinsic logarithmic scale of EFC. Integrating this expression gives:
$U(\vec{Z})=\dfrac{1}{2}\sum_{i,j}A_{ij}e^{Z_{i}}e^{Z_{j}}-\sum_{i}Z_{i}+\delta\sum_{i}e^{Z_{i}}$ (20)
Finally, Eq. (7) can be obtained by replacing the original variables $V_{i}=e^{-Z_{i}}$.
Appendix B NHEFC: Solution in the case N=2
In the case of a simple network consisting of $N=2$ connected nodes, the equations defining the iterative map can be directly formulated. Specifically, we have
$\begin{cases}V^{(n+1)}_{1}=\delta+\frac{1}{V^{(n)}_{2}},\\ V^{(n+1)}_{2}=\delta+\frac{1}{V^{(n)}_{1}}\end{cases}$
along with the associated function
$U(\vec{V})=\frac{1}{V_{1}V_{2}}+\log V_{1}+\log V_{2}+\frac{\delta}{V_{1}}+\frac{\delta}{V_{2}}.$
From these expressions, it follows that
$V^{(n+2)}_{j}=f(V^{(n)}_{j}),\quad j=1,2,\quad\text{where }f(x)=\delta+\frac{x}{1+\delta x}.$
The iterative dynamics given by $x_{n+1}=f(x_{n})$ can be analyzed using a Lamerey diagram. Observing that the function $f(x)$ has a unique fixed point in the interval $[0,+\infty)$, given by
$x^{*}=\frac{1}{2}\left(\sqrt{4+\delta^{2}}+\delta\right),$
and that its derivative satisfies
$0<f^{\prime}(x)<1,\quad f([0,+\infty))=[\delta,1/\delta),$
it follows that the sequence $x_{n}$ converges monotonically to $x^{*}$: if $x_{0}\in[0,x^{*})$, then $x_{n}$ is increasing and approaches $x^{*}$, while if $x_{0}>x^{*}$, the sequence is decreasing and converges to $x^{*}$.
Furthermore, the function $U(\vec{V})$ reaches a stationary point within the domain $[0,+\infty)^{2}$ if and only if $\vec{V}=(x^{*},x^{*})$, as determined by the condition $\nabla U(\vec{V})=0$. Since $U(\vec{V})\to+\infty$ when $\vec{V}$ approaches the boundary of $[0,+\infty)^{2}$, this unique stationary point is necessarily the absolute minimizer of $U$. Consequently, both sequences $U(\vec{V}^{(2n)})$ and $U(\vec{V}^{(2n+1)})$ converge to $\min U$, ensuring that the entire sequence $U(\vec{V}^{(n)})$ also converges to its minimum value.
It is important to emphasize the critical role of the positive parameter $\delta$ in ensuring convergence toward the fixed point. In particular, when $\delta=0$, there remains a unique fixed point, given by $\vec{V}^{*}=(1,1)$. However, in this case, since the map effectively alternates between two steps at a time and reduces to the identity function, $f(x)=x$, the system does not exhibit true convergence. Instead, the phase space becomes entirely populated by periodic orbits, each with a period of two, oscillating around $\vec{V}^{*}$.
From a dynamical perspective, the parameter $\delta$ plays a role analogous to that of linear friction in a mechanical system. Just as friction causes the minimum of a potential function to transition from being merely stable to asymptotically stable, the presence of $\delta>0$ ensures that trajectories are progressively damped and eventually settle at the fixed point, eliminating the periodic behavior observed in the frictionless case.
Appendix C NHEFC: Regular graphs
The iterative scheme is defined by
$V^{(n)}_{i}=\delta+\sum_{j}\frac{A_{ij}}{V^{(n-1)}_{j}},$
where the matrix elements $A_{ij}$ are such that, for some integer $K\in\{1,\ldots,N-1\}$,
$K_{i}:=\sum_{j}A_{ij}=K\qquad\forall\,i.$
In particular, if $K=N-1$, the graph is complete (clique). Notice that, independently of the initial condition $\vec{V}^{(0)}\in(0,+\infty)^{N}$, the iterates are uniformly bounded. One easily verifies that
$\delta\leq V^{(n)}_{i}\leq\delta+\frac{K}{\delta}\qquad\forall\,n\geq 1\quad\forall\,i.$ (21)
To analyze the convergence properties of the map, we define the minimum and maximum values at iteration $n$ as
$V^{(n)}_{m}:=\min_{j}V^{(n)}_{j},\qquad V^{(n)}_{M}:=\max_{j}V^{(n)}_{j}.$
Then, by the structure of the update, it follows that
$\delta+\frac{K}{V^{(n)}_{M}}\leq V^{(n+1)}_{i}\leq\delta+\frac{K}{V^{(n)}_{m}}\qquad\forall\,i.$
This immediately implies the bounds
$\delta+\frac{K}{V^{(n)}_{M}}\leq V^{(n+1)}_{m}\quad\text{and}\quad V^{(n+1)}_{M}\leq\delta+\frac{K}{V^{(n)}_{m}}.$
By iterating these inequalities once more, we obtain
$V^{(n+2)}_{m}\geq f_{K}(V^{(n)}_{m}),\qquad V^{(n+2)}_{M}\leq f_{K}(V^{(n)}_{M}),$ (22)
where we have introduced the function
$f_{K}(x)=\delta+\frac{Kx}{K+\delta x}.$
It is instructive to note that
$0<f_{K}^{\prime}(x)<1,\quad f_{K}([0,+\infty))=[\delta,K/\delta),$
which implies that the iterative process
$X^{(n+1)}=f_{K}(X^{(n)})$ (23)
behaves in a manner analogous to the two-node case. In particular, for any initial condition $X^{(0)}>0$, the sequence $X^{(n)}$ converges to the unique fixed point $V^{*}$ of $f_{K}$ in $[0,\infty)$, where
$V^{*}=\sqrt{K+\tfrac{1}{4}\delta^{2}}+\delta.$
To further formalize this convergence, denote by $X^{(n)}_{m}$ and $Y^{(n)}_{m}$ the orbits starting from initial values $X^{(0)}_{m}\leq V^{(0)}_{m}$ and $Y^{(0)}_{m}\leq V^{(1)}_{m}$, respectively. Then, by the inequalities in (22) and the monotonicity of $f_{K}$,
$V^{(2n)}_{m}\geq X^{(n)}_{m},\qquad V^{(2n+1)}_{m}\geq Y^{(n)}_{m}\qquad\forall\,n\geq 0.$
These bounds can be rigorously established via induction. For instance, assuming that $V^{(2n)}_{m}\geq X^{(n)}_{m}$, we have
$V^{(2n+2)}_{m}\geq f_{K}(V^{(2n)}_{m})\geq f_{K}(X^{(n)}_{m})=X^{(n+1)}_{m}.$
A similar argument, using orbits $X^{(n)}_{M}$ and $Y^{(n)}_{M}$ starting from $X^{(0)}_{M}\geq V^{(0)}_{M}$ and $Y^{(0)}_{M}\geq V^{(1)}_{M}$, leads to
$V^{(2n)}_{M}\leq X^{(n)}_{M},\qquad V^{(2n+1)}_{M}\leq Y^{(n)}_{M}\qquad\forall\,n\geq 0.$
In summary, the sequences $\{V^{(n)}_{m}\}$ and $\{V^{(n)}_{M}\}$ are bounded respectively from below and above by trajectories of the dynamical system defined in (23). Consequently, we obtain
$\lim_{n\to\infty}\vec{V}^{(n)}=\vec{V}^{*}\qquad\forall\,\vec{V}^{(0)}\in(0,+\infty)^{N},$ (24)
with $\vec{V}^{*}:=(V^{*},\ldots,V^{*})$.
It is worthwhile to note that, from (22) and the lower bound in (21), one deduces that
$V^{(n)}_{M}-V^{(n)}_{m}\leq\frac{K^{2}\left(V^{(n-2)}_{M}-V^{(n-2)}_{m}\right)}{(\delta V^{(n-2)}_{M}+K)(\delta V^{(n-2)}_{m}+K)}\leq\frac{K^{2}}{(K+\delta^{2})^{2}}\left(V^{(n-2)}_{M}-V^{(n-2)}_{m}\right).$
Since $\frac{K^{2}}{(K+\delta^{2})^{2}}<1$, this inequality shows that the convergence in (24) is exponentially fast in time. Moreover, applying (24) to the stationary case, we deduce that the fixed point equation
$V_{i}=\delta+\sum_{j}\frac{A_{ij}}{V_{j}}$ (25)
admits a unique solution $\vec{V}=\vec{V}^{*}$ in $(0,+\infty)^{N}$.
As discussed at the end of Subsection 3.2.2 in the explicit case $N=2$, the presence of the positive parameter $\delta$ is crucial for ensuring convergence to a unique fixed point. When $N>2$ and $\delta=0$, the argument leading to (24) fails because the function $f_{K}(x)$ becomes the identity $f_{K}(x)=x$. Moreover, unlike the special case $N=2$, it is not even clear whether $\vec{V}^{*}=(\sqrt{K},\ldots,\sqrt{K})$ is the unique fixed point.
Regarding the relation to the cost function, we observe that
$\frac{\partial U}{\partial V_{i}}=-\frac{1}{V_{i}^{2}}\sum_{j}\frac{A_{ij}}{V_{j}}+\frac{1}{V_{i}}-\frac{\delta}{V_{i}^{2}}=-\frac{1}{V_{i}^{2}}\left(\sum_{j}\frac{A_{ij}}{V_{j}}-V_{i}+\delta\right).$
Hence, the condition for a stationary point, $\nabla U(\vec{V})=0$, is equivalent to the fixed point equation (25). Since $U(\vec{V})$ diverges to +$\infty$+ ∞ as $\vec{V}$ approaches the boundary of $(0,+\infty)^{N}$, we conclude that the fixed point $\vec{V}^{*}$ is indeed the unique absolute minimizer of $U$ in $(0,+\infty)^{N}$.
Appendix D NHEFC: local convergence for general graphs
We write the dynamics in vector form as
$\vec{V}^{(n+1)}=\vec{F}(\vec{V}^{(n)})$
where $\vec{F}(\vec{V})$ denotes the vector-valued function with components
$\vec{F}_{i}(\vec{V})=\delta+\sum_{j}\frac{A_{ij}}{V_{j}}$
Drawing upon the results established in Section 3.2.1, the mapping $\vec{F}$ possesses precisely one fixed point, which coincides with the unique minimizer $\vec{V}^{*}$ of the cost function $U$. We aim to demonstrate that $W=U-\min U$ serves as a Lyapunov function for the equilibrium $\vec{V}^{*}$. According to standard stability theory, this would imply that the equilibrium is asymptotically stable.
Evidently, there exists a ball $B=B_{\varepsilon}(\vec{V}^{*})$ of radius $\varepsilon$ centered at $\vec{V}^{*}$ within which $W\geq 0$ and $W(\vec{V})=0$ if and only if $\vec{V}=\vec{V}^{*}$. To complete the proof, we must show that, for sufficiently small $\varepsilon$,
$U(F(\vec{V}))-U(\vec{V})<0\qquad\forall\,\vec{V}\in B\setminus\{\vec{V}^{*}\}$
(which implies that the same property holds with $W$ in place of $U$). To this end, we observe that
$F_{i}(\vec{V})=V_{i}-V_{i}^{2}\frac{\partial U}{\partial V_{i}}(\vec{V}),$
Therefore, defining $\vec{D}=\vec{F}(\vec{V})-\vec{V}$, we find
$\nabla U(\vec{V})\cdot\vec{D}=-\sum_{i}\frac{D_{i}^{2}}{V_{i}^{2}}$
By Taylor’s formula,
$\begin{split}U(F(\vec{V}))-U(\vec{V})&=\nabla U(\vec{V})\cdot\vec{D}+\int_{0}^{1}\!\mathrm{d}\lambda\,\lambda\vec{D}\cdot D^{2}U(\vec{V}+\lambda\vec{D})\vec{D}\\ &=-\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}+\frac{1}{2}\vec{D}\cdot D^{2}U(\vec{V}^{*})\vec{D}+\mathcal{O}(\varepsilon)|\vec{D}|^{2}\qquad\forall\,\vec{V}\in B,\end{split}$
where in the last equality, we have utilized the fact that if $\vec{V}\in B$, then $|\vec{V}+\lambda\vec{D}-\vec{V}^{*}|\leq C\varepsilon$, since $\vec{F}$ is regular and $\vec{V}^{*}$ is a fixed point. Furthermore,
$(D^{2}U)_{ij}(\vec{V})=\frac{A_{ij}}{V_{i}^{2}V_{j}^{2}}+\left[\frac{2}{V_{i}^{3}}\left(\sum_{s}\frac{A_{ij}}{V_{s}}+\delta\right)-\frac{1}{V_{i}^{2}}\right]\delta_{ij}$
The Hessian evaluated at the fixed point simplifies to
$(D^{2}U)_{ij}(\vec{V}^{*})=\frac{A_{ij}}{(V_{i}^{*})^{2}(V_{j}^{*})^{2}}+\frac{\delta_{ij}}{(V_{i}^{*})^{2}}$
(note that it is positive definite, by the same argument used for Eq. (9)), yielding
$\frac{1}{2}\vec{D}\cdot D^{2}U(\vec{V}^{*})\vec{D}=\frac{1}{2}\sum_{i,j}\frac{A_{ij}D_{i}D_{j}}{(V_{i}^{*})^{2}(V_{j}^{*})^{2}}+\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}$
Applying the Cauchy–Schwarz inequality and leveraging the fixed point property of $\vec{V}^{*}$, we obtain
$\begin{split}\frac{1}{2}\sum_{i,j}\frac{A_{ij}D_{i}D_{j}}{(V_{i}^{*})^{2}(V_{j}^{*})^{2}}&=\frac{1}{2}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\frac{D_{i}}{V_{i}^{*}}\frac{D_{j}}{V_{j}^{*}}\leq\frac{1}{4}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\left(\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}+\frac{D_{j}^{2}}{(V_{j}^{*})^{2}}\right)\\ &=\frac{1}{2}\sum_{i,j}\frac{A_{ij}}{V_{i}^{*}V_{j}^{*}}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}\\ &=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}\sum_{j}\frac{A_{ij}}{V_{j}^{*}}=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}(V_{i}^{*}-\delta)\\ &=\frac{1}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{2}}-\frac{\delta}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}\end{split}$
In conclusion, denoting $V_{M}^{*}=\max_{i}V_{i}^{*}$, we establish
$U(F(\vec{V}))-U(\vec{V})\leq-\frac{\delta}{2}\sum_{i}\frac{D_{i}^{2}}{(V_{i}^{*})^{3}}+\mathcal{O}(\varepsilon)|\vec{D}|^{2}\leq\left(-\frac{\delta}{2(V_{M}^{*})^{3}}+\mathcal{O}(\varepsilon)\right)|\vec{D}|^{2}\qquad\forall\,\vec{V}\in B$
which implies $U(F(\vec{V}))-U(\vec{V})<0$ provided $\varepsilon$ is sufficiently small and $\vec{D}\neq\vec{0}$ (i.e., $\vec{V}\neq\vec{V}^{*}$).